OpenStudy (anonymous):

how to differentiate (x-y)^4

OpenStudy (anonymous):

\[4(x-y)^3 \times \left(1 - \frac{dy}{dx}\right)\]

OpenStudy (anonymous):

\[x^4 - y^4\] is this right?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Aww -.- i really need to work my simple math.